OpenStudy (anonymous):

Heather paid $1,940.20 for a computer. If the price paid includes a 9% sales tax, which of the following equations can be used to determine the price of the computer after tax? (Let x represent the cost of the computer and y represent the total cost after tax.) A. y = 1.9x B. y = x + 9x C. y = 0.91x D. y = 1.09x 6. Marcus works as a salesman at a car dealership. He is paid a base salary of $1,253.20 each month, and he receives a commission of $161.45 for each vehicle he sells. If last month Marcus earned $8,195.55, how many cars did he sell last month? A. 29 B. 43 C.

OpenStudy (anonymous):

A. 29 B. 43 C. 86 D. 102

OpenStudy (anonymous):

#5) Let "x' be whatever item's price is and 9% could be presented as 0.09 or 9/100. but we want that tax to apply to the item's price. so we need it to be 1 + 0.09, where "1" is the initial value of the price of the item. #6) 8195.55 = 1253.20 + 161.45c where "c" represents the number of cars sold last month. As 1253.20 is his base salary for each month and the total earnings he got is 8195.55

OpenStudy (anonymous):

So #5 = and D?

OpenStudy (anonymous):

Before, I answer, why would you choose that as your answer?

OpenStudy (anonymous):

Well, I thought that you would combine the "1" and 0.9 so it would make 1.09 as tax.

OpenStudy (anonymous):

True. But what's so special about the "1" why is that "1" there?

OpenStudy (anonymous):

well 1 can for now be the initial value of the price item. Like you said.

OpenStudy (anonymous):

Right. So say the price of chair would be $55.00 and the sales tax is 9%. Say we do y = (55.00)(0.09) = $4.95, which is way below our initial value. that's why we have the "1" there. y = (55.00)(1.09) = $59.95

OpenStudy (anonymous):

now solve for #6 and explain why.

OpenStudy (anonymous):

So wait the answer for #5 is?

OpenStudy (anonymous):

A. y = 1.9x B. y = x + 9x C. y = 0.91x D. y = 1.09x

OpenStudy (anonymous):

Lol. Why are you unsure? You have given me the answer and I asked why you got that answer =)

OpenStudy (anonymous):

Oh D! :D

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Okay so #6 8195.55 = 1253.20 + 161.45c fill in c c = 86 Check: 8195.55 = 1253.20 + 161.45(86)

OpenStudy (anonymous):

Okay. Good job. But do you know how I came up with that equation?

OpenStudy (anonymous):

Nope! And thats what were learning in class how to come up with them.

OpenStudy (anonymous):

Okay well I was incorrect. The answer is not "c" for #6. Seems like you did your math incorrectly. So it told us that Marcus made a total of $8,195.55 last month right? and then they told us that he gets $1,253.20 every month, and lastly he gets $161.45 for each car sold. And we're asked to find the amount of cards sold, that is why I put that "c" there

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

So I set $8,195.55 as the "value" we want to get to as that's what Marcus got last month. and then set $1,253.20 as our initial value since Marcus gets that every month right? then our dependent value $161.45 relied on the amount of cars sold in which if you read and understood, this equation is represented: $8,195.55 = $1,253.20 + $161.45(amount of cars 'c' )

OpenStudy (anonymous):

okay,

OpenStudy (anonymous):

So what is the answer for #6?

OpenStudy (anonymous):

same thing?

OpenStudy (anonymous):

thats the same equation.

OpenStudy (anonymous):

Yeah. But I told you did your math incorrectly. Look at it closely and do it again.

OpenStudy (anonymous):

Okay!

OpenStudy (anonymous):

B!

OpenStudy (anonymous):

AWESOME! Alright. Cheers to you.

OpenStudy (anonymous):

Thank you so much! :D

OpenStudy (anonymous):

You're very welcome.

OpenStudy (anonymous):

:D